The width of a rectangular box is 4 times its height, and its length is 3 ft more than its height. The volume of the box is 49 ft cube. Find the height of the box.
volume = lwh
49 = (4x)(3+x)(x)
49 = 4x3 +12x2

Answer by stanbon(48545)

The width of a rectangular box is 4 times its height, and its length is 3 ft more than its height. The volume of the box is 49 ft cube. Find the height of the box.
volume = lwh
------------------------
Let the height be "x":
width = 4x
length = x+3
-----------------------
EQUATION:
V = lwh
49 = x*4x*(x+3)
49 = 4x^3 + 12x^2
4x^3 + 12x^2 - 49 = 0
I graphed it and found:
x = 1.627098...ft (height
4x = 6.508392
x+3= 4.627
============
Cheers,
Stan H.